OpenStudy (anonymous):

Determine the slope and y-intercept of the line whose equation is given : 7y+12x-8=0

OpenStudy (kc_kennylau):

It would be easier if you could convert the equation into slope-intercept form :)

OpenStudy (anonymous):

the problem is ı can't. in the textbook's answer key it says : y=12/6x-4/6 but how did it happen?

OpenStudy (kc_kennylau):

Do you notice anything special about the left hand side of any equation in slope-intercept form?

OpenStudy (kc_kennylau):

It only contains "y" :)

OpenStudy (kc_kennylau):

Therefore, to convert any equation to slope-intercept form, you just have to make "y" the subject :)

OpenStudy (kc_kennylau):

To do so, you need to throw all things that do NOT contain "y" to the right hand side.

OpenStudy (anonymous):

actually ı understood that part, but ı think the main problem in here is where did "12/6 and 4/5 came from"?

OpenStudy (kc_kennylau):

Your textbook is wrong

OpenStudy (kc_kennylau):

The correct answer is completely different from the answer in the book

OpenStudy (anonymous):

what is your answer?

OpenStudy (kc_kennylau):

y=-12/7 x + 8/7

OpenStudy (anonymous):

for 2y plus 5x-8=0 it says y=1/2x plus 2 it says that. is that wrong too?

OpenStudy (kc_kennylau):

exactly

OpenStudy (kc_kennylau):

However, I don't mean not to trust you, but I doubt that you're intentionally posting the wrong answers from the textbook in order that we do your homework for you. Sorry for any mistunderstanding.

OpenStudy (kc_kennylau):

sorry for not trusting you

OpenStudy (anonymous):

actually ı just wanted to learn how can ı do that. because there are two similar questions and ı don't seem to understand why ı can't do that

OpenStudy (kc_kennylau):

\[Ax+By+C=0\]\[By=-Ax-C\]\[y=\frac{-Ax-C}B\]

OpenStudy (kc_kennylau):

As I said, throw all the things not related to "y" to the right hand side first, then you can use multiplication or division to clear the rest of the stuffs sticking to "y".

OpenStudy (anonymous):

thank you so much.

OpenStudy (kc_kennylau):

no problem at all :)